write an equation in standard form with roots -4,10
Answer by nerdybill(7384)

For a quadratic to have -4,10 as roots then it must have the following factors:
(x+4)(x-10)
FOIL the above:
x^2 -10x + 4x - 40
x^2 -6x - 40
.
putting it in standard form:
x^2 -6x - 40=0
(x^2 -6x +9) = 40 + 9
(x-3)^2 = 49
f(x) = (x-3)^2 -49